John sold 5 rolls of plain wrapping paper and 8 rolls of shiny wrapping paper for a total of $199. Totsaka sold 5 rolls of plain wrapping paper and 6 rolls of shiny wrapping paper for a total of $163.
Determine the cost each of one roll of plain wrapping paper and one roll of shiny wrapping paper?
-Thanks to anyone who helps.

Answer by mananth(16946)

x cost of roll of plain wrapping paper
y cost of roll of shiny paper
Total attended
5 x + 8 y = 199 .............1

5 x + 6 y = 163 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 4
-15 x -24 y = -597
20 x + 24 y = 652
Add the two equations
5 x = 55
/ 5
x = 11
plug value of x in (1)
5 x + 8 y = 199
55 + 8 y = 199
8 y = 199 -55.00
8 y = 144
y = 18
$ 11 cost of roll of plain wrapping paper
$ 18 cost of roll of shiny paper
m.ananth@hotmail.ca